# Evolving Symbolic 3D Visual Grounder with Weakly Supervised Reflection

### v2

@@ Abstract
3D visual grounding (3DVG) is challenging because of the requirement of understanding on visual information, language and spatial relationships. While supervised approaches have achieved superior performance, they are constrained by the scarcity and high cost of 3D vision-language datasets. On the other hand, LLM/VLM based agents are proposed for 3DVG, eliminating the need for training data. However, these methods incur prohibitive time and token costs during inference. To address the challenges, we introduce a novel training-free symbolic framework for 3D visual grounding, namely Evolvable Symbolic Visual Grounder (EaSe ), that offers significantly reduced inference costs compared to previous agent-based methods while maintaining comparable performance. EaSe uses LLM generated codes to compute on spatial relationships. EaSe also implements an automatic pipeline to evaluate and optimize the quality of these codes and integrate VLMs to assist in the grounding process. Experimental results demonstrate that EaSe achieves 52.9% accuracy on Nr3D dataset and 49.2% Acc@0.25 on ScanRefer, which is top-tier among training-free methods. Moreover, it substantially reduces the inference time and cost, offering a balanced trade-off between performance and efficiency. Codes are available at https://github.com/OpenRobotLab/EaSe.
@@
The 3D visual grounding (3DVG) task focuses on locating an objects in a 3D scene based on a referring utterance. Numerous supervised methods have been proposed for 3DVG (Hsu et al., 2023; Jain et al., 2022; Huang et al., 2022; Chen et al., 2022; Huang et al., 2024; Ziyu et al., 2023; Bakr et al., 2023; Wu et al., 2023). These methods learn representations of referring utterance, object attributes and spatial relations from large-scale 3D vision language training data with high-quality annotations and achieve SOTA performances on 3DVG. However, 3D vision language datasets (Chen et al., 2020; Achlioptas et al., 2020) are scarce and their annotations are expensive. Additionally, some supervised methods are trained on these close-vocabulary datasets, which restricts their applicability in open-vocabulary scenarios. In recent years, large language models (LLMs) and vision-language models (VLMs) have shown remarkable capabilities in reasoning, code generation and visual perception. Building on these advancements, open-vocabulary and zero shot 3DVG agents (Yang et al., 2023; Xu et al., ; Fang et al., 2024; Li et al., 2024a) are proposed. These methods let LLM perform numerical computing and reasoning on object locations in the text modality (Yang et al., 2023; Fang et al., 2024), or let a VLM locate targets from scene scan images in the visual modality (Xu et al., ). Their responses which contain planning, reasoning or self-debugging processes are usually long. However, the generation is online, which means for every grounding query, the LLM generates long responses. As a result, they have significant costs in terms of grounding time and token usage (see Agents block in LABEL:fig:banner).
In contrast, visual programming method (Yuan et al., 2024b) utilizes LLM to generate a program which uses annotated relation functions and outputs the target object by executing the program. The generated program is short so its time and token consumption is much lower. However, it has trouble considering many spatial relations in the referring utterance simultaneously (Yuan et al., 2024a). This causes its relatively low accuracy (see Visprog. in LABEL:fig:banner).
To the end, we introduce EaSe, a novel training-free symbolic framework integrates LLM and VLM for 3D visual grounding which balances accuracy and inference cost. EaSe derives from previous neuro-symbolic frameworks (Hsu et al., 2023; Feng et al., 2024) and use LLM generated and optimized Python codes as spatial relation encoders (see  Figure 1, block (b)). EaSe also employs a VLM to process visual information of objects, which is also important but can’t be handled by code.
Specifically, EaSe parses the referring utterance into a symbolic expression which encapsulates all mentioned object categories and their spatial relations. Given positions of all objects in the scene, the spatial relation encoders generate relation features which can represent spatial relations between them. Then a special executor aggregates the symbolic expression, relation features and object categories to exclude most of objects that don’t match the referring utterance. Finally, the VLM selects the target from images of remaining candidate objects (see Figure 1, block (c)).
For more accurate relation encoders, we generate them through iterative optimization processes instead of straightly prompting the LLM. We introduce test suites to evaluate relation encoders. Not only the test suites enable us to choose better relation encoders from LLM responses, but also the LLM can leverage failed test cases to optimize its responses. Compared to the online generation of agent based methods, our generation and optimization is offline. Our relation encoders are reused in all grounding processes, so EaSe has advantages on time and token cost (see Ours in LABEL:fig:banner).
@@
We evaluate EaSe on the widely used ScanRefer (Chen et al., 2020) and Nr3D (Achlioptas et al., 2020) datasets. Experimental results show that EaSe  achieves 52.9% accuracy on Nr3D and 49.2% Acc@0.25 on ScanRefer, matching the performance of agent-based approaches (Xu et al., ; Fang et al., 2024) while offering significant advantages in grounding time and token cost. In paticular, on Nr3D, EaSe  outperforms VLMGrounder (Xu et al., ) in accuracy while requiring less than $1/20$ of the grounding time and less than $1/6$ of the token usage. In addition, EaSe  significantly outperforms ZSVG3D (Yuan et al., 2024b) in accuracy with comparable grounding time and token cost. In conclusion, among various training-free 3DVG methods, EaSe  strikes an excellent balance between accuracy and efficiency.
@@
Figure 1: Overview of EaSe. Block (a) and (b) compare the difference of relation encoders between EaSe and previous neuro-symbolic approaches. Relation encoders compute relation features using object positions from the scene. The executor executes the symbolic expression with relation features and gets candidate objects. Block (c) demonstrates a VLM selects the target from candidate objects given scan images containing them.
2 2. Related Works
@@
3D visual grounding focuses on finding an object in 3D scene based on descriptions to appearance or location. ScanRefer (Chen et al., 2020) and ReferIt3D (Achlioptas et al., 2020) are two popular benchmarks on this task, providing rich and diverse object-utterance pairs on ScanNet (Dai et al., 2017). Many supervised approaches train a end-to-end model for 3DVG. Jain et al. (2022) combines bottom-up and top-down object detection and grounding with transformer. Chen et al. (2022) designs fine-grained neural spatial relation encoders. Ziyu et al. (2023) achieves promising performance by training on large scale well annotated dataset. Compared to supervised approaches, EaSe doesn’t rely on large scale training data. Neuro-symbolic approaches (Hsu et al., 2023; Feng et al., 2024; Li et al., 2024b) use LLMs to parse the referring utterance into symbolic expressions and learnable neural networks as spatial relation encoders. In contrast, EaSe use Python codes are spatial relation encoders.
Yuan et al. (2024b) manually annotates functions and prompt LLM to generate function calling code. Yang et al. (2024); Fang et al. (2024) create agents where LLM or VLM can reason and take actions iteratively. Xu et al. uses VLM and images from the scene to figure out the target object without the need for detection or segmentation modules. Inference cost of EaSe is as low as Yuan et al. (2024b) and the performance is comparable as Yang et al. (2024); Xu et al. ; Fang et al. (2024).
@@
LLM can generate executable codes (Roziere et al., 2023; Zhu et al., 2024) from scratch or according to in context documents.  Li et al. (2023a) executes generated codes for precise mathematical reasoning.  Liang et al. (2023) uses LLM to generate policy codes for robot controlling.  Xie et al. ; Ma et al. (2023) prompt LLM to design reward functions for RL training and use feedback from real world to improve them. Zhou et al. (2024) generates functions to refine pre-training corpora. Gupta and Kembhavi (2023); Yuan et al. (2024b) annotate tool APIs and use LLM to convert natural language instructions into API-calling codes. LLMs are also able to refine its generated code based on error message from executor or feedback from environment. Le et al. (2022); Chen et al. (2023) use the feedback from the outside environment to fine-tuning language models, Ma et al. (2023) uses RL training trajectories to select and improve its reward functions. In 3DVG area, Yuan et al. (2024b); Fang et al. (2024) also use code to process spatial relations, But to the best of our knowledge, we are the first to design specific test suites for code selection and improvement.
Neuro-symbolic Reasoning
Neuro-symbolic reasoning methods parse natural language to symbolic expression and get reasoning results by executing it with neural networks. Li et al. (2023b) provides a neuro-symbolic interface that supports many logical reasoning and seamless integration with other pre-trained model seamlessly. Cheng et al. (2023) uses LLM query to assist SQL execution. Mao et al. (2019); Hsu et al. (2023); Feng et al. (2024) define domain-specific language and use neural networks as relation encoders for 3DVG. EaSe uses similar symbolic frameworks, but our relation encoders are Python codes. Spatial knowledge of our encoders is from LLM instead of 3D training data.
@@ 3.1 Problem Statement
3D visual grounding tasks involve a scene, denoted as $\mathcal{S}$, represented by an RGB-colored point cloud containing $C$ points, where $\mathcal{S}\in\mathbb{R}^{C\times 6}$. Accompanying this is an utterance $\mathcal{U}$ that describes an object within the scene $\mathcal{S}$. The objective is to identify the location of the target object $\mathcal{T}$ in the form of a 3D bounding box. In the ReferIt3D dataset (Achlioptas et al., 2020), bounding boxes for all objects are provided, making the visual grounding process a task of matching these bounding boxes to the scene $\mathcal{S}$. In contrast, the ScanRefer dataset (Chen et al., 2020) provides only the scene point cloud, requiring additional detection or segmentation modules to accomplish the grounding task.
@@ 3.2 Grounding Pipeline
We adhere to the previous SOTA neuro-symbolic framework for 3DVG (Hsu et al., 2023; Feng et al., 2024). A semantic parser (LLM) converts $\mathcal{U}$ into a symbolic expression $\mathcal{E}$ in JSON. Relation encoders compute the features for relations such as under and right within $\mathcal{E}$. Finally, the features, along with category features of relevant object classes in $\mathcal{E}$ (“chair" and “desk"). These features and logits are subsequently used to calculate the matching degrees between $\mathcal{S}$ and the objects based on $\mathcal{E}$.
@@
We employ GPT-4o (OpenAI, 2024) to parse $\mathcal{U}$ into JSON expression $\mathcal{E}$, which contains the categories and spatial relations in $\mathcal{U}$. For example, the utterance “chair near the table" can be represented as:
@@
Human-annotated natural language expressions exhibit diverse descriptions of relations, leading to a long-tail distribution of relation_name in parsed expressions. To mitigate this, we predefine a set of common relation names and prompt LLM to select from them for $\mathcal{E}$ instead of using the original terms from $\mathcal{U}$.
Based on the number of associated objects, the relations can be categorized into unary, binary, and ternary (Feng et al., 2024). For simplicity, attributes that describe properties of a single object, such as “large" or “at the corner" are treated as special types of unary relations. We presents our predefined set of relations along with their classifications in Table 1.
@@
Neuro-symbolic approaches(Hsu et al., 2023; Feng et al., 2024) uses neural network as relation encoders. In contrast, our relation encoders are Python codes generated by LLMs. It take the scene’s point cloud data and object segmentation as input and compute spatial relation features of objects in an explicit way. The object category features are from a pre-trained point cloud classifier. The feature of unary relations and category features, denoted as $f_{\text{unary}}\in\mathbb{R}^{N}$, can be seen as the similarity between objects and their respective categories or relations, where $N$ is the number of objects in the scene. Features of binary relations $f_{\text{binary}}\in\mathbb{R}^{N\times N}$ represent the likelihood of binary relations existing between all possible pairs of objects. For instance, the element $f_{\text{near}}^{i,j}$ quantifies the probability that the $i$-th object is “near" the $j$-th object. The structure of ternary features follows a similar pattern, capturing relations involving three objects.
Execution
Our executor is similar to Hsu et al. (2023). The executor utilizes the symbolic expression $\mathcal{E}$ and the associated relation features to identify the target object $\mathcal{T}$. Since the elements in the features represent probabilities or corresponding relations/categories, logical conjunctions within $\mathcal{E}$ are modeled using product operations.
For a given symbolic expression $\mathcal{E}$: The executor use $\mathcal{E}$ and associated features to identify $\mathcal{T}$. Since elements in the features represent the probabilities or corresponding relation or category, the logical conjunction in $\mathcal{E}$ can be represented through the product operation. For a symbolic expression $\mathcal{E}$, its category feature $f_{\text{category}}$ of calculated initially by a classifier. Subsequently, the executor processes each relation individually of the relations field. For the $i$-th relation, the relation feature $f_{\text{relation}}$ is computed and multiplied with $f_{\text{category}}$ of its related objects, resulting a $f_{\text{i}}\in\mathbb{R}^{N}$. After processing all relations, we obtain multiple relation features, each of size $\mathbb{R}^{N}$. They are aggregated with $f_{\text{category}}$ through an element-wise product to compute the final matching degree.
3.3 Relation Encoders
The sizes and positions of objects in 3D scenes are intrinsically linked to specific spatial relations. For instance, “near” depends on the distance between objects, while “large” depends on their volume. In EaSe , each of these spatial relations is handled by a dedicated piece of Python code that computes its associated feature.
The generation and optimization of these codes are illustrated in Figure 2. During the initial code generation stage, an optimized snippet for the relevant relation is retrieved as an in-context example (§$\S$ 3.3.2). Then we sample multiple responses from LLM and test them with a test suite (§$\S$ 3.3.3). Any failures in the test suite trigger error message synthesis, guiding the LLM to iteratively refine and optimize the codes (§$\S$ 3.3.4). The following sections provide a more detailed explanation of this process.

### v3

@@ Abstract
3D visual grounding (3DVG) is challenging due to the need to understand 3D spatial relations. While supervised approaches have achieved superior performance, they are constrained by the scarcity and high cost of 3D vision-language datasets. Training-free approaches based on LLM/VLM agents eliminate the need for training data, but they incur prohibitive grounding time and token costs. To address the challenges, we introduce a novel training-free symbolic framework for 3D visual grounding, namely Evolvable Symbolic Visual Grounder (EaSe ). EaSe uses LLM-generated code to determine 3D spatial relations among objects and integrates VLMs to process their visual information. EaSe also implements an automatic pipeline that evaluates and optimizes the quality of generated code. Experimental results demonstrate that EaSe achieves 52.9% accuracy on Nr3D dataset and 49.2% Acc@0.25 on ScanRefer, ranking among the best training-free methods. Moreover, it substantially reduces the grounding time and token cost, offering a balanced trade-off between performance and efficiency. Code is available at https://github.com/OpenRobotLab/EaSe.
@@
The 3D visual grounding (3DVG) task focuses on locating an object in a 3D scene based on a referring utterance. Numerous supervised methods have been proposed for 3DVG (Hsu et al., 2023; Jain et al., 2022; Huang et al., 2022; Chen et al., 2022; Huang et al., 2024; Ziyu et al., 2023; Bakr et al., 2023; Wu et al., 2023). These methods learn representations of referring utterances, object attributes and spatial relations from large scale 3D vision-language training data with high-quality annotations and achieve state-of-the-art performances on 3DVG. However, the scarcity of 3D vision-language datasets (Chen et al., 2020; Achlioptas et al., 2020), coupled with the high cost of their annotations, limits their applicability. Furthermore, some supervised methods are trained on these closed-vocabulary datasets, restricting their applicability in open-vocabulary scenarios.
In recent years, large language models (LLMs) and vision-language models (VLMs) have shown remarkable capabilities in reasoning, code generation, and visual perception. Building on these advancements, open-vocabulary and zero-shot 3DVG agents (Yang et al., 2023; Xu et al., ; Fang et al., 2024; Li et al., 2024a) are proposed. These methods let LLMs perform numerical computing and reasoning on object locations in the text modality (Yang et al., 2023; Fang et al., 2024), or let VLMs locate targets from scene scan images in the visual modality (Xu et al., ). Leveraging the reasoning and perceptual abilities of advanced LLMs and VLMs, these agents achieve superior accuracy compared to other training-free methods. However, they rely on LLMs to produce lengthy responses (containing planning, reasoning, or self-debugging processes) for every referring utterance. This online generation style results in significant costs in terms of grounding time and token usage (see Agents block in LABEL:fig:banner). In contrast, the visual programming method (Yuan et al., 2024b) utilizes LLM to generate a program which uses annotated relation functions and outputs the target object by executing the program. The generated program is short so its time and token consumption are much lower. However, it has trouble considering many spatial relations in the referring utterance simultaneously (Yuan et al., 2024a). This results in relatively low accuracy (see Visprog. in LABEL:fig:banner).
To address these dual challenges of accuracy and cost, we propose EaSe, a novel training-free symbolic framework that integrates LLMs and VLMs for 3D visual grounding, balancing both accuracy and inference cost. EaSe builds upon previous neuro-symbolic frameworks (Hsu et al., 2023; Feng et al., 2024) and uses Python code generated and optimized by LLMs as spatial relation encoders (see  Figure 1, block (b)). EaSe also employs a VLM to distinguish objects that differ only in visual appearance. Specifically, EaSe parses the referring utterance into a symbolic expression which encapsulates all mentioned object categories and their spatial relations. Given positions of all objects in the scene, the spatial relation encoders generate relation features which can represent spatial relations between them. Then an executor aggregates the symbolic expression, relation features and object categories to exclude most objects that do not match the referring utterance. Finally, the VLM identifies the target object from images of the remaining candidate objects (see Figure 1, block (c)). For more accurate spatial relation encoders, we generate them through iterative optimization processes instead of directly prompting the LLM. We introduce test suites to evaluate spatial relation encoders. The test suites not only enable us to select better relation encoders from LLM responses but also allow the LLM to leverage failed test cases to optimize its code. In contrast to the online generation of agent-based methods, our generation and optimization are performed offline, avoiding per-utterance code generation. The spatial relation encoders are reused in all grounding processes, so EaSe has advantages on time and token cost (see Ours in LABEL:fig:banner).
@@
We evaluate EaSe on the widely used ScanRefer (Chen et al., 2020) and Nr3D (Achlioptas et al., 2020) datasets. Experiment results show that EaSe  achieves 52.9% accuracy on Nr3D and 49.2% Acc@0.25 on ScanRefer, matching the performance of agent-based approaches (Xu et al., ; Fang et al., 2024) while offering significant advantages in grounding time and token cost. On Nr3D, EaSe  outperforms VLMGrounder (Xu et al., ) in accuracy while requiring less than $1/20$ of the grounding time and less than $1/6$ of the token usage. In addition, EaSe  significantly outperforms ZSVG3D (Yuan et al., 2024b) in accuracy with comparable grounding time and token cost. In conclusion, among various training-free 3DVG methods, EaSe  strikes an excellent balance between accuracy and efficiency.
@@
Figure 1: Overview of EaSe. Block (a) and (b) compare the difference of spatial relation encoders between EaSe and previous neuro-symbolic approaches. Spatial relation encoders compute relation features using object positions from the scene. The executor executes the symbolic expression with relation features and gets candidate objects. Block (c) demonstrates a VLM selects the target from candidate objects given scan images containing them.
2 2. Related Work
@@
3D visual grounding (3DVG) aims to localize objects in 3D scenes based on natural language descriptions of appearance and spatial relations. Two dominant benchmarks, ScanRefer (Chen et al., 2020) and ReferIt3D (Achlioptas et al., 2020), leverage ScanNet (Dai et al., 2017) scenes to provide diverse object-utterance pairs. Supervised Methods typically train end-to-end models on annotated 3D vision-language data. Jain et al. (2022) integrates bottom-up object detection with transformer-based grounding, Chen et al. (2022) designs fine-grained neural networks to encode spatial relations. Though achieving promising accuracy, these methods suffer from expensive data annotation dependency (Ziyu et al., 2023). Neuro-Symbolic Methods (Hsu et al., 2023; Feng et al., 2024; Li et al., 2024b) attempt to mitigate data reliance by combining symbolic parsing with neural components. They parse referring utterances into symbolic expressions using LLMs and train neural networks as spatial relation encoders. Unlike these approaches, EaSe completely avoids training on large-scale 3D datasets. Training-free methods exploit pre-trained LLMs / VLMs for open-vocabulary 3DVG. Yuan et al. (2024b) uses LLMs to generate programs that call predefined functions to find the target object. Yang et al. (2024); Fang et al. (2024) deploy LLM/VLM-based agents that analyze object appearances and locations and find the target. Xu et al. uses VLMs and images from the scene to figure out the target object. Concurrently, Yuan et al. (2024a) proposes to replace the programming of Yuan et al. (2024b) by constraint satisfaction solving. Li et al. (2024a) parses landmark and perspective of the referring utterance and then uses VLMs to find the target object from a rendered image. Compared to these methods, EaSe offers a superior balance of accuracy and efficiency, outperforming  Yang et al. (2024); Xu et al. ; Fang et al. (2024) in grounding time and token cost while surpassing Yuan et al. (2024b) in accuracy.
@@
LLMs demonstrate growing proficiency in generating executable code (Roziere et al., 2023) for precise mathematical reasoning (Li et al., 2023), robotics control (Liang et al., 2023), tool use (Gupta and Kembhavi, 2023; Yuan et al., 2024b) or data cleaning (Zhou et al., 2024). Recent work further explores code refinement via environmental feedback, such as RL training trajectories (Ma et al., 2023) or real-world execution errors (Le et al., 2022; Chen et al., 2023). In the 3DVG area, Yuan et al. (2024b); Fang et al. (2024) also uses code to process spatial relations, but EaSe advances this paradigm by introducing test suites to automatically optimize code.
@@ 3.1 Problem Statement
3D visual grounding tasks involve a scene, denoted as $\mathcal{S}$, represented by an RGB-colored point cloud containing $C$ points, where $\mathcal{S}\in\mathbb{R}^{C\times 6}$. Associated with this is an utterance $\mathcal{U}$ that describes an object within the scene $\mathcal{S}$. The objective is to identify the location of the target object $\mathcal{T}$ in the form of a 3D bounding box. In the ReferIt3D dataset (Achlioptas et al., 2020), bounding boxes for all objects are provided, making the visual grounding process a task of matching these bounding boxes to the scene $\mathcal{S}$. In contrast, the ScanRefer dataset (Chen et al., 2020) provides only the point cloud of the scene, requiring additional detection or segmentation modules to accomplish the grounding task.
@@ 3.2 Grounding Pipeline
We adhere to the previous SOTA neuro-symbolic framework for 3DVG (Hsu et al., 2023; Feng et al., 2024). A semantic parser (LLM) converts $\mathcal{U}$ into a symbolic expression $\mathcal{E}$ in JSON. Spatial relation encoders compute the relation features such as under and right within $\mathcal{E}$. Finally, the relation features, along with category features of object categories in $\mathcal{E}$, such as chair and desk are subsequently used to calculate the matching scores between $\mathcal{S}$ and the objects based on $\mathcal{E}$.
@@
We employ GPT-4o (OpenAI, 2024) to parse $\mathcal{U}$ into JSON expression $\mathcal{E}$, which contains the categories and spatial relations in $\mathcal{U}$. For example, the utterance “chair near the table” can be represented as:
@@
Human-annotated natural language expressions exhibit diverse descriptions of relations, leading to a long-tail distribution of relation_name in parsed expressions. To mitigate this, we define a set of common relation names and prompt LLM to select from them for $\mathcal{E}$ instead of using the original terms from $\mathcal{U}$.
Based on the number of associated objects, the relations can be categorized into unary, binary, and ternary (Feng et al., 2024). For simplicity, attributes that describe properties of a single object, such as “large” or “at the corner” are treated as special types of unary relations. We present our predefined set of relations along with their classifications in Table 1.
@@
Our spatial relation encoders are Python code generated by LLMs. They take the objects’ point clouds and positions as input and compute spatial relation features by explicit geometric calculations. The category features are from a pretrained point cloud classifier. Unary relation features $f_{\text{unary}}\in\mathbb{R}^{N}$ and category features $f_{\text{category}}\in\mathbb{R}^{N}$ measure the similarity between objects and their respective relations or categories, where $N$ is the number of objects in the scene. The features of the binary relation $f_{\text{binary}}\in\mathbb{R}^{N\times N}$ represent the likelihood that there are binary relations between all possible pairs of objects. For example, $f_{\text{near}}^{(i,j)}$ quantifies the probability that the $i$-th object is near the $j$-th object. Ternary features follow an analogous pattern for relations involving three objects.
Executor
Our executor follows a similar design to Hsu et al. (2023). Given the symbolic expression $\mathcal{E}$ and features, the executor computes the matching score between objects and referring utterance $\mathcal{U}$. Specifically, the category features $f_{\text{category}}$ are computed by an object classifier, and relation features are from our spatial relation encoders. For each relation in $\mathcal{E}$’s relations field, the corresponding relation feature $f_{\text{relation}}$ is multiplied with $f_{\text{category}}$ of its related objects, yielding intermediate features $\{f_{i}\in\mathbb{R}^{N}\}_{i=1}^{K}$ (where $K$ is the number of relations). Finally, all intermediate features and $f_{\text{category}}$ are aggregated via the element-wise product to compute the final matching score. See Algorithm 2 for more details.
3.3 Spatial Relation Encoders
Sizes and positions of objects in 3D scenes inherently determine spatial relations. For example, the near relation depends on pairwise distances, while large is determined by object volumes. In EaSe , each spatial relation is handled by a dedicated Python class (§$\S$ 3.3.1) that can compute its associated features given the object bounding boxes.
As illustrated in Figure 2, our encoder generation involves three phases: (1) Retrieving in-context examples from semantically similar relations (§$\S$ 3.3.2), (2) Sampling multiple code implementations from LLMs, and (3) Validating candidates through unit tests (§$\S$ 3.3.3). When test failures occur, we automatically synthesize error messages that guide iterative refinement (§$\S$ 3.3.4).